# Permutation groups, partition lattices and block structures

### v2

@@ 4.2.1 General results
For example, the transitive actions of $S_{5}$ and $A_{5}$ on $15$ points are both quasiprimitive but not pre-primitive. However, there is a unique non-trivial invariant partition in each case, with $5$ parts each of size $3$; so, by Corollary 4.4, these groups are OB.
@@
Proposition 4.7
The OB property is preserved by direct and wreath product.
This follows from the argument on page 2.4.
@@ 4.2.2 Products
First we prove some general facts about invariant partitions of direct products of an arbitrary number of groups in their product action, and slightly extend a result in [1], proving that the direct product of an arbitrary number of primitive groups in its product action is pre-primitive. This result is interesting in its own right, but it will also be used to show that a generalised wreath product of primitive groups is pre-primitive, a key fact that we will use in the proof of our main theorem. First we give some language to describe partitions of products.
@@
[…]
  • Choose a fixed $\delta\in\Delta$, and consider the intersections of the parts of $\Pi$ with $\Gamma\times\{\delta\}$. These form a partition of $\Gamma\times\{\delta\}$ and so, by ignoring the second factor, we obtain a partition of $\Gamma$ called the $G$-fibre partition. Now the action of the group $\{1\}\times H$ shows that it is independent of the element $\delta\in\Delta$ chosen.
@@
First we observe that the projection and fibre partitions on $\Gamma$ agree if and only if those on $\Delta$ agree. For the pairs in a part $P$ of $\Pi$ are the edges of a bipartite graph on $A\cup B$, where $A$ and $B$ are parts of the projection partitions; the valency of a point in $A$ is equal to the number of points of $B$ in a part of the fibre partition on $\Delta$, which we will denote by $a$; and similarly the valencies $b$ of the points in $B$. Then counting edges of the graph (that is, pairs in part $P$ of $\Pi$), we see that $|A|a=|B|b$. Now the fibre and projection partitions on $\Gamma$ agree if and only if $|A|=b$, which is equivalent to $|B|=a$.
@@
Next we introduce the notion of partition orthogonality. Let $G,H$ be transitive permutation groups, on $\Gamma$, $\Delta$ respectively, as above. We say that $G$ and $H$ are partition-orthogonal if the only $G\times H$-invariant partitions of $\Gamma\times\Delta$ are of the form $\{\Gamma_{i}\times\Delta_{j}\mid i\in\{1,\ldots,m\},j\in\{1,\ldots,n\}\}$ where $\{\Gamma_{1},\ldots,\Gamma_{m}\}$ is a $G$-invariant partition of $\Gamma$ and $\{\Delta_{1},\ldots,\Delta_{n}\}$ is an $H$-invariant partition of $\Delta$.
@@
We prove the claim by induction. If $m=2$, then the claim follows by the definition. Suppose that the claim holds for $m-1$ factors. Let $H=G_{1}\times\cdots\times G_{m-1}$ and suppose for a contradiction that there is some $G$-invariant partition $\Pi$ which is not a direct product of partitions of the sets $\Omega_{i}$. Then the $H$-fibre and $H$-projection partitions induced on $\Omega_{1}\times\cdots\times\Omega_{m-1}$ by $\Pi$ must differ.
@@
We prove the claim by induction on $m$.
We first prove the claim for $m=2$. Suppose for a contradiction that there exists a $(G_{1}\times G_{2}\times H)$-invariant partition $\Pi$ which is not the result of crossing a $(G_{1}\times G_{2})$-invariant partition of $\Omega_{1}\times\Omega_{2}$ with an $H$-invariant partition of $\Delta$. Thus, there exist $(\alpha_{1},\beta_{1},\delta_{1})$ and $(\alpha_{2},\beta_{2},\delta_{2})$ in the same part of $\Pi$ with $(\alpha_{1},\beta_{1})\neq(\alpha_{2},\beta_{2})$, and for these pairs, necessarily $\delta_{1}\neq\delta_{2}$. This implies that the projection and fibre partitions of $\Pi$ onto either $\Omega_{1}\times\Delta$ or $\Omega_{2}\times\Delta$ must differ, contradicting the assumption that $G_{i}$ and $H$ are partition-orthogonal for the relevant $i\in\{1,2\}$.
@@
Let $\Pi$ be a $G\times H$-invariant partition of $\Gamma\times\Delta$. Since $G$ and $H$ are partition-orthogonal, $\Pi$ is the direct product of a $G$-invariant partition $\Pi_{G}$ and an $H$-invariant partition $\Pi_{H}$. Since both $G$ and $H$ are pre-primitive, it follows that $\Pi_{G}$ and $\Pi_{H}$ are orbit partitions of some subgroups $M$ and $N$ of $G$ and $H$ respectively. It is then easy to check that $\Pi$ is the orbit partition of $M\times N$, which proves the claim.   $\Box$
@@ 4.2.3 Regular groups
It follows from Corollary 4.2 that, if $G$ is a regular permutation group, then $G$ has the OB property if and only if any two subgroups of $G$ commute. These groups were determined by Iwasawa [23]; we refer to Schmidt [29, Chapter 2] for all the material we require. In this section we use the term quasi-hamiltonian, taken from [14], for a group in which any two subgroups commute. (The term will not be used outside this section.)
@@
We have not found a reference for precisely what we want, so we give a direct proof of the first part; the second is [29, Theorem 2.3.1].
@@
Note that not every quasi-hamiltonian group is a Dedekind group; so the OB property lies strictly between transitivity and pre-primitivity. Note also that $Q_{8}$ is quasi-hamiltonian but $Q_{8}\times Q_{8}$ is not; so the OB property is not closed under direct product.
@@
We first show that the subgroups containing $G_{\alpha}$ are of the form $HG_{\alpha}$ for some $H\leq N$ invariant under the action of $G_{\alpha}$. Let $M$ be such a subgroup. Since $M\leq G=NG_{\alpha}$ all the elements of $M$ are of the form $ng$ where $n\in N$ and $g\in G_{\alpha}$. Then since $G_{\alpha}\leq M$ it follows that $n=ngg^{-1}\in M$. Hence, $M=HG_{\alpha}$, where $H=N\cap M\leq N$.
@@ 4.2.4 Modularity and distributivity
We have seen that the subgroup lattice of a group, which clearly determines modularity, does not determine whether the subgroups commute. So we cannot expect a characterisation of the OB property in terms of the lattice of subgroups containing a given point stabiliser. But is there anything to say here?
@@
This is true because a group with distributive subgroup lattice is locally cyclic, by Ore’s theorem [29, Section 1.2], and a finite locally cyclic group is cyclic. Since a cyclic group is Dedekind, it is pre-primitive and so has the OB property.
@@ 4.3.1 A group-theoretic result
Let $G$ be the generalised wreath product of the groups $G(m)$ over a poset $M$, acting on a set $\Omega$. Let $p$ be a minimal element of $M$. Let $\Pi$ be the corresponding partition of $\Omega$, $H$ the group induced on the set of parts by $G$, $N$ the stabiliser of all parts of $\Pi$. Then
@@
    $N$ is a direct product of copies of $G(p)$, where there is an equivalence relation $\sim$ on the parts of $\Pi$ (determined by the poset $M$) such that each direct factor acts in the same way on the parts in one equivalence class and fixes every point in the other parts;
@@
Part (b) is proved using the definition of a generalised wreath product. The equivalence relation is defined as follows: for parts $\mu$ and $\sigma$ of $\Pi$, $\mu\sim\sigma$ if and only if $\mu$ and $\sigma$ lie in the same part of $\Pi\vee\Phi$ for all partitions $\Phi$ of the poset block structure defined by $M$ which are incomparable to $\Pi$.
First note that since $N$ fixes the parts of $\Pi$, it must also fix the parts of every partition lying above $\Pi$. Therefore, only parts of partitions incomparable to $\Pi$ can be moved by $N$. Now let $\Phi$ denote a partition incomparable to $\Pi$. Note that since $\Pi\preccurlyeq\Pi\vee\Phi$, the parts of $\Phi$ contained in the same part of $\Pi\vee\Phi$ can only be permuted amongst themselves by $N$. Hence, if $h\in N$ simultaneously acts in the same way on two parts, say $\mu,\sigma$ of $\Pi$, then those two parts must be contained in the same part of $\Pi\vee\Phi$ for every partition $\Phi$ incomparable to $\Pi$.
It now remains to show that if $\mu,\sigma\in\Pi$ are such that $\mu\sim\sigma$, then $N$ acts in the same way on $\mu$ and $\sigma$. Let $\gamma,\delta$ lie in $\mu$ and $\sigma$, and moreover suppose that they are contained in the same part of $\Phi$ for every partition $\Phi$ incomparable to $\Pi$. It suffices to show that every $h\in N$ maps $\gamma$ and $\delta$ to the same part of $\Phi$ for every $\Phi$ incomparable to $\Pi$.
Now $h$ can be written as a product $\prod_{\Phi}h_{\Phi}$, where each factor $h_{\Phi}$ encodes the permutation of the parts of the corresponding partition $\Phi$ induced by $h$. Hence, it suffices to show that $h_{\Phi}$ maps $\gamma$ and $\delta$ to the same part for an arbitrary $\Phi$ incomparable to $\Pi$. We may assume without loss of generality that $\Phi$ is join-indecomposable, since every element is a join of JI elements, and the distributive law implies that if a collection of JI elements are incomparable with $\Pi$ then so is their join.
@@
We finally note that $\sim$ is only dependent on the poset $M$ and not the group $G$.
For (c), we have to show that $H$ normalises $N$ and that the action of $H$ extends to $\Omega$. The first statement is clear since $N$ is the subgroup fixing all parts of $\Pi$. For the second, note that $H$ acts on the set of $(|M|-1)$-tuples; extend each element to act on $|M|$-tuples by acting as the identity on the $p$-th coordinate.  $\Box$
@@ 4.3.2 Generalised wreath products of primitive groups
In this section, we will use the notation for poset block structures and generalised wreath products defined in Section 2.5. Moreover, let $[N]$ denote the set $\{1,\ldots,N\}$, and for every subset $J$ of $M$, let $X_{J}$ be the index set of $J$, namely $\{i\in[N]\,:\,m_{i}\in J\}$. We then define $P_{J}$ to be the partition whose set of parts is
@@
Let $K$ denote the direct product of the groups $G(m_{i})$ (for $i\in[N]$) in its product action and $P$ denote the lattice of partitions corresponding to down-sets in $M$.
@@
for all $i\in[N]$, and therefore, $fh\in H$. We now show that the function $z=\prod_{i\in[N]}z_{i}$ where $z_{i}$ sends all elements of $\Omega^{i}$ to the inverse of $\mathop{\mathrm{im}}(f_{i})$ is the inverse of $f$ and therefore that $f^{-1}\in H$. Indeed,
@@
(b) First note that if the only partitions preserved by $G$ are the ones in $P$, then clearly $G$ is PB. Suppose first that there are further partitions other than the ones corresponding to down-sets in $M$ fixed by $G$, and let $\Pi$ be such a partition. Since $K\leq G$, it follows that $\Pi$ is also preserved by $K$. Therefore, $\Pi$ must have one of the two following forms:
  • $\Pi$ is of the form
@@
  • at least two of the $G(m_{i})$s, say $G(m_{i})$ and $G(m_{j})$, are cyclic of the same prime order, and $\Pi$ is a partition whose corresponding $G(m_{i})$ and $G(m_{j})$-fibre partitions are the partitions into singletons and the $G(m_{i})$ and $G(m_{j})$-projection partitions are those with a single part.
@@
Hence, $\Pi$ must be of the second type and thus there exist $G(m_{i})$ and $G(m_{j})$ cyclic of prime order and $\Pi$ is a partition whose corresponding $G(m_{i})$ and $G(m_{j})$-fibre partitions are the partitions into singletons and the $G(m_{i})$ and $G(m_{j})$-projection partitions are the partitions into a single part. If $m_{i}$ and $m_{j}$ are related, say $m_{i}\sqsubset m_{j}$ then, as above, $\Pi$ must be preserved by the iterated wreath product $G(a_{0})\wr G(a_{1})\wr\ldots\wr G(a_{k})$. However, knowing what partitions imprimitive iterated wreath products preserve, we deduce that $G(a_{0})\wr G(a_{1})\wr\ldots\wr G(a_{k})$ cannot preserve $\Pi$ and therefore $m_{i}$ and $m_{j}$ must be incomparable.
Now we have to show the converse. So suppose that there are two incomparable elements in $M$, say $m_{1}$ and $m_{2}$, such that the corresponding groups are cyclic of the same prime order $p$. As defined in Section 2.5,

### v3

@@ 4.2.1 General results
For example, the transitive actions of $S_{5}$ and $A_{5}$ on $15$ points (on the cosets of a Sylow 2-subgroup) are both quasiprimitive but not pre-primitive. However, there is a unique non-trivial invariant partition in each case, with $5$ parts each of size $3$; so, by Corollary 4.4, these groups are OB.
@@ 4.2.2 Products
First we prove some general facts about invariant partitions of direct products of an arbitrary number of groups in their product action, and slightly extend a result in [1], proving that the direct product of an arbitrary number of primitive groups in its product action is pre-primitive. This result is interesting in its own right, but it will also be used to show that a generalised wreath product of primitive groups is pre-primitive, which constitutes a part of Theorem 4.19. First we give some language to describe partitions of products.
@@
    We claim that the sets $P_{0}$ arising in this way are pairwise disjoint. For suppose that $\gamma\in P_{0}\cap Q_{0}$, where $Q_{0}$ is defined similarly for another part $Q$ of $\Pi$; suppose that $(\gamma,\delta_{1})\in P$ and $(\gamma,\delta_{2})\in Q$. There is an element $h\in H$ mapping $\delta_{1}$ to $\delta_{2}$. Then $(1,h)$ maps $(\gamma,\delta_{1})$ to $(\gamma,\delta_{2})$, and hence maps $P$ to $Q$, and $P_{0}$ to $Q_{0}$; but this element acts trivially on $\Gamma$, so $P_{0}=Q_{0}$. It follows that the sets $P_{0}$ arising in this way form a partition of $\Gamma$, which we call the $G$-projection partition of $\Gamma$ induced by $\Pi$.
  • Choose a fixed $\delta\in\Delta$, and consider the intersections of the parts of $\Pi$ with $\Gamma\times\{\delta\}$. These form a partition of $\Gamma\times\{\delta\}$ and so, by ignoring the second factor, we obtain a partition of $\Gamma$ called the $G$-fibre partition of $\Gamma$ induced by $\Pi$. Now the action of the group $\{1\}\times H$ shows that it is independent of the element $\delta\in\Delta$ chosen.
@@
First we observe that the projection and fibre partitions on $\Gamma$ agree if and only if those on $\Delta$ agree. For the pairs in a part $P$ of $\Pi$ are the edges of a bipartite graph on $A\cup B$, where $A$ and $B$ are parts of the projection partitions on $\Gamma$ and $\Delta$ respectively; the valency of a point in $A$ is equal to the number of points of $B$ in a part of the fibre partition on $\Delta$, which we will denote by $a$; and similarly the valencies $b$ of the points in $B$. Then counting edges of the graph (that is, pairs in part $P$ of $\Pi$), we see that $|A|a=|B|b$. Now the fibre and projection partitions on $\Gamma$ agree if and only if $|A|=b$, which is equivalent to $|B|=a$.
@@
Next we introduce the notion of partition orthogonality.
Definition
Let $G,H$ be transitive permutation groups, on $\Gamma$, $\Delta$ respectively, as above. We say that $G$ and $H$ are partition-orthogonal if the only $G\times H$-invariant partitions of $\Gamma\times\Delta$ are of the form $\{\Gamma_{i}\times\Delta_{j}\mid i\in\{1,\ldots,m\},j\in\{1,\ldots,n\}\}$ where $\{\Gamma_{1},\ldots,\Gamma_{m}\}$ is a $G$-invariant partition of $\Gamma$ and $\{\Delta_{1},\ldots,\Delta_{n}\}$ is an $H$-invariant partition of $\Delta$.
@@
We prove the claim by induction. If $m=2$, then the claim follows by the definition. Suppose that the claim holds for m1m-1 factors. Let $H=G_{1}\times\cdots\times G_{m-1}$ and suppose for a contradiction that there is some $G$-invariant partition $\Pi$ which is not a direct product of partitions of the sets $\Omega_{i}$. Then the $H$-fibre and $H$-projection partitions induced on $\Omega_{1}\times\cdots\times\Omega_{m-1}$ by $\Pi$ must differ.
@@
Let $G=G_{1}\times\cdots\times G_{m}\times H$ and $\Omega=\Omega_{1}\times\cdots\times\Omega_{m}\times\Delta$. We prove the claim by induction on $m$.
We first prove the claim for $m=2$. Let $(\alpha,\beta,\delta),(\alpha^{\prime},\beta^{\prime},\delta^{\prime})\in\Omega_{1}\times\Omega_{2}\times\Delta$. First consider the $H$-fibre partition $\Pi_{\Delta}$ of $\Delta$ induced by $\Pi$. Note that by definition $(\alpha,\beta,\delta)$ and $(\alpha,\beta,\delta^{\prime})$ are in the same part of $\Pi$ if and only if $\delta$ and $\delta^{\prime}$ are in the same part of $\Pi_{\Delta}$.
Now consider the $(G_{1}\times H)$-fibre partition $\Pi_{\Omega_{1}\times\Delta}$ of $\Omega_{1}\times\Delta$ induced by $\Pi$. Since $G_{1}$ and $H$ are partition-orthogonal by assumption, we must have $\Pi_{\Omega_{1}\times\Delta}=\Pi_{\Omega_{1}}\times\Pi_{\Delta}$, where $\Pi_{\Omega_{1}}$ denotes the $G_{1}$-fibre partition of $\Omega_{1}$ induced by $\Pi$. This means that $(\alpha,\beta,\delta)$ and $(\alpha^{\prime},\beta,\delta^{\prime})$ are in the same part of $\Pi$ if and only if $\alpha$ and $\alpha^{\prime}$ are in the same part of $\Pi_{\Omega_{1}}$ and $\delta$ and $\delta^{\prime}$ are in the same part of $\Pi_{\Delta}$.
An entirely similar argument shows that $(\alpha,\beta,\delta)$ and $(\alpha,\beta^{\prime},\delta^{\prime})$ are in the same part of $\Pi$ if and only if $\beta$ and $\beta^{\prime}$ are in the same part of the $G_{2}$-fibre partition $\Pi_{\Omega_{2}}$ of $\Omega_{2}$ induced by $\Pi$ and $\delta$ and $\delta^{\prime}$ are in the same part of $\Pi_{\Delta}$.
It therefore follows that $(\alpha,\beta,\delta)$ and $(\alpha^{\prime},\beta^{\prime},\delta^{\prime})$ are in the same part of $\Pi$ if and only if $\alpha$ and $\alpha^{\prime}$ are in the same part of $\Pi_{\Omega_{1}}$, $\beta$ and $\beta^{\prime}$ are in the same part of $\Pi_{\Omega_{2}}$ and $\delta$ and $\delta^{\prime}$ are in the same part of $\Pi_{\Delta}$, and so
$\Pi=\Pi_{\Omega_{1}}\times\Pi_{\Omega_{2}}\times\Pi_{\Delta}$
which proves the claim.
@@
Let $\Pi$ be a $G\times H$-invariant partition of $\Gamma\times\Delta$. Since $G$ and $H$ are partition-orthogonal, $\Pi$ is the direct product of a $G$-invariant partition $\Pi_{G}$ and an $H$-invariant partition $\Pi_{H}$. Since both $G$ and $H$ are pre-primitive, it follows that $\Pi_{G}$ and $\Pi_{H}$ are orbit partitions of some subgroups $G^{*}$ and $H^{*}$ of $G$ and $H$ respectively. It is then easy to check that $\Pi$ is the orbit partition of $G^{*}\times H^{*}$, which proves the claim.   $\Box$
@@ 4.2.3 Regular groups
It follows from Corollary 4.2 that, if $G$ is a regular permutation group, then $G$ has the OB property if and only if any two subgroups of $G$ commute. These groups were determined by Iwasawa [25]; we refer to Schmidt [31, Chapter 2] for all the material we require. In this section we use the term quasi-hamiltonian, taken from [15], for a group in which any two subgroups commute. (The term will not be used outside this section.)
@@
We have not found a reference for precisely what we want, so we give a direct proof of the first part; the second is [31, Theorem 2.3.1].
@@
Note that not every quasi-hamiltonian group is a Dedekind group, namely a group all of whose subgroups are normal; so the OB property lies strictly between transitivity and pre-primitivity. Note also that $Q_{8}$ (acting regularly) is quasi-hamiltonian but $Q_{8}\times Q_{8}$ in the product action is not; so the OB property is not closed under direct product.
@@
We first show that the subgroups containing $G_{\alpha}$ are of the form $HG_{\alpha}$ for some $H\leq N$ invariant under the action of $G_{\alpha}$. Let $F$ be a subgroup containing $G_{\alpha}$. Since $F\leq G=NG_{\alpha}$ all the elements of $F$ are of the form $ng$ where $n\in N$ and $g\in G_{\alpha}$. Then since $G_{\alpha}\leq F$ it follows that $n=ngg^{-1}\in F$. Hence, $F=HG_{\alpha}$, where $H=N\cap F\leq N$. Since both $F$ and $N$ are invariant under the action of $G_{\alpha}$, so is $H$.
@@ 4.2.4 Modularity and distributivity
We have seen at the start of Section 4.2.3 that the subgroup lattice of a group, which clearly determines modularity, does not determine whether the subgroups commute. So we cannot expect a characterisation of the OB property in terms of the lattice of subgroups containing a given point stabiliser. But is there anything to say here?
@@
This is true because a group with distributive subgroup lattice is locally cyclic, by Ore’s theorem [31, Section 1.2], and a finite locally cyclic group is cyclic. Since a cyclic group is Dedekind, it is pre-primitive and so has the OB property.
@@ 4.3.1 A group-theoretic result
Let $G$ be the generalised wreath product of the groups $G(m)$ over a poset $M$, acting on a set $\Omega$. Let $p$ be a minimal element of $M$. Let $\Pi$ be the corresponding partition of $\Omega$, $H$ the group induced on the set of parts by $G$, NN the stabiliser of all parts of $\Pi$. Then
@@
    NN is a direct product of copies of $G(p)$, where there is an equivalence relation $\sim$ on the set of parts of $\Pi$ (determined by the poset $M$) such that each direct factor acts in the same way on the parts in one equivalence class and fixes every point in the other parts;
@@
Part (b) is proved using the definition of a generalised wreath product. The equivalence relation is defined as follows: for parts $P$ and $Q$ of $\Pi$, $P\sim Q$ if and only if $P$ and $Q$ lie in the same part of $\Pi\vee\Phi$ for all partitions $\Phi$ of the poset block structure defined by $M$ which are incomparable to $\Pi$.
First note that since $N$ fixes the parts of $\Pi$, it must also fix the parts of every partition lying above $\Pi$. Therefore, only parts of partitions incomparable to $\Pi$ can be moved by $N$. Now let $\Phi$ denote a partition incomparable to $\Pi$. Note that since $\Pi\preccurlyeq\Pi\vee\Phi$, the parts of $\Phi$ contained in the same part of $\Pi\vee\Phi$ can only be permuted amongst themselves by $N$. Hence, if the actions of $h\in N$ on two parts of $\Pi$ are equivalent, then those two parts must be contained in the same part of $\Pi\vee\Phi$ for every partition $\Phi$ of $\Omega$ incomparable to $\Pi$.
It now remains to show that if $P,Q\in\Pi$ are such that $P\sim Q$, then $N$ acts in the same way on $P$ and $Q$. Let $\gamma,\delta$ lie in $P$ and $Q$, and moreover suppose that they are contained in the same part of $\Phi$ for every partition $\Phi$ of $\Omega$ incomparable to $\Pi$. It suffices to show that every $h\in N$ maps $\gamma$ and $\delta$ to the same part of $\Phi$ for every $\Phi$ incomparable to $\Pi$.
Now $h$ can be written as a product $\prod_{\Phi}h_{\Phi}$, where each factor $h_{\Phi}$ encodes the permutation induced by $h$ of the parts of the corresponding partition $\Phi$ of $\Omega$ induced by $h$. Hence, it suffices to show that $h_{\Phi}$ maps $\gamma$ and $\delta$ to the same part for an arbitrary partition $\Phi$ of $\Omega$ incomparable to $\Pi$. We may assume without loss of generality that $\Phi$ is join-indecomposable, since every element is a join of JI elements, and the distributive law implies that if a collection of JI elements are incomparable with $\Pi$ then so is their join.
@@
We finally note that $\sim$ is dependent only on the poset $M$ and not the group $G$.
For (c), we have to show that $H$ normalises $N$ and that the action of $H$ extends to $\Omega$. The first statement is clear since $N$ is the subgroup fixing all parts of $\Pi$. For the second, note that $H$ acts on the set of (|M|1)(|M|-1)-tuples; extend each element to act on $|M|$-tuples by acting as the identity on the $p$-th coordinate.  $\Box$
@@ 4.3.2 Generalised wreath products of primitive groups
In this section, we will use the notation for poset block structures and generalised wreath products defined in Section 2.5. Moreover, let $[N]$ denote the set $\{1,\ldots,N\}$, and for every subset $J$ of $M$, let $X_{J}$ be the index set of $J$, namely $\{i\in[N]\,:\,m_{i}\in J\}$. We then define $P_{J}$ to be the partition of $\Omega$ whose set of parts is
@@
Let $K$ denote the direct product of the groups $G(m_{i})$ (for $i\in[N]$) in its product action and PP denote the lattice of partitions corresponding to down-sets in $M$.
@@
for all $i\in[N]$, and therefore, $fh\in H$. For $i$ in $[N]$, let $z_{i}$ be the function which maps all the elements of $\Omega^{i}$ to the inverse of ${\rm im}(f_{i})$. Put $z=\prod_{i\in[N]}z_{i}$. Then
@@
We now prove (b).
(ii) $\Rightarrow$ (i). Note that if the only partitions preserved by $G$ are the ones in $P$, then clearly $G$ is PB.
(iii) $\Rightarrow$ (ii). We prove the contrapositive. Suppose that there are further partitions fixed by $G$ other than the ones corresponding to down-sets in $M$, and let $\Pi$ be such a partition. Since $K\leq G$, it follows that $\Pi$ is also preserved by $K$. Therefore, by the arguments in the proof of Theorem 4.13 we deduce that since all of the $G(m_{i})$s are primitive, one of the following must hold:
  • All the $G(m_{i})$s are mutually partition-orthogonal, and so $\Pi$ is of the form
@@
  • at least two of the $G(m_{i})$s, say $G(m_{i})$ and $G(m_{j})$, are cyclic of the same prime order, and $\Pi$ is a partition whose corresponding $G(m_{i})$ and $G(m_{j})$-fibre partitions are different to the $G(m_{i})$ and $G(m_{j})$-projection partitions respectively. Since the degree of $G(m_{i})$ and $G(m_{j})$ is prime, this can only happen if the fibre partitions are the partitions into singletons and the projection partitions are the partitions into a single part.
@@
Hence, $\Pi$ must be of the second type and thus there exist $G(m_{i})$ and $G(m_{j})$ cyclic of the same prime order and $\Pi$ is a partition whose corresponding $G(m_{i})$ and $G(m_{j})$-fibre partitions are the partitions into singletons and the $G(m_{i})$ and $G(m_{j})$-projection partitions are the partitions into a single part. If $m_{i}$ and $m_{j}$ are related, say $m_{i}\sqsubset m_{j}$ then, as above, $\Pi$ must be preserved by the iterated wreath product $G(a_{0})\wr G(a_{1})\wr\ldots\wr G(a_{k})$. However, knowing what partitions imprimitive iterated wreath products preserve, we deduce that $G(a_{0})\wr G(a_{1})\wr\ldots\wr G(a_{k})$ cannot preserve $\Pi$ and therefore $m_{i}$ and $m_{j}$ must be incomparable.
(i) $\Rightarrow$ (iii). We again prove the contrapositive. So suppose that there are two incomparable elements in $M$, say $m_{1}$ and $m_{2}$, such that the corresponding groups are cyclic of the same prime order $p$. As defined in Section 2.5,